# Nearly-polynomial inverse theorem for the $U^{d}$ norm in degree $d+1$

### 17 Mar 2026 v1

@@ 6.2 Main proof
We now put everything together and prove Theorem 1.
@@
$\operatorname{rk}(g)\leq\operatorname{rk}(\overline{g})\quad\text{ and }\quad\operatorname{rk}(h-\partial_{c}g)\leq\operatorname{rk}(\overline{h}-\partial_{c}\overline{g}).$ (7)
To see this, abbreviate $\mathbf{x}^{t}=(x,\ldots,x)$ ($t$ times), and observe that $\partial_{c}\overline{g}(\mathbf{x}^{d})=\overline{\partial_{c}g}(\mathbf{x}^{d})=d!\partial_{c}g(x)$, using Lemma 5.2 and Fact 5.7. Moreover by Fact 5.7, we have $\overline{h}(\mathbf{x}^{d})=d!h(x)$ and $\overline{g}(\mathbf{x}^{d+1})=(d+1)!g(x)$. As $\operatorname{ch}(\mathbb{F})>d+1$, we thus have $g(x)=\overline{g}(\mathbf{x}^{d})/(d+1)!$ and $(h-\partial_{c}g)(x)=(\overline{h}-\partial_{c}\overline{g})(\mathbf{x}^{d})/d!$, proving (7).
By Proposition 6.4, $\operatorname{bias}(\Delta^{d}f)\leq\operatorname{bias}(\overline{g})$, and $\operatorname{bias}(\Delta^{d}f)\leq\operatorname{bias}(\overline{h}-\partial_{c}\overline{g})$ for some direction vector $c$ which we henceforth fix. Put $\operatorname{bias}(\Delta^{d}f)=q^{-r}$, so that
@@
Now, combine (7), (8), and Theorem 2.1 to deduce that
@@
$\operatorname{rk}^{*}(g+h)\leq 2\operatorname{rk}(g)+\operatorname{rk}(h-\partial_{c}g)\leq\tilde{O}_{d}(r).$
Next, apply Proposition 4.7 with $k=d+1$ ($\geq 3$ by assumption) on $g+h\in\operatorname{Poly}_{d+1}$. We deduce that there exists a polynomial $P$ with $\deg(P)\leq k-2=d-1$ such that
$\operatorname{cor}(g+h,\,P)\geq q^{-2\operatorname{rk}^{*}(g+h)}\geq 1/q^{\tilde{O}_{d}(r)}.$
To finish, write $f=g+h+f_{<d-1}$. We claim that $P+f_{<d-1}\in\operatorname{Poly}_{d-1}$ significantly correlates with $f$; indeed,
$\operatorname{cor}_{<d}(f)$ $\geq\operatorname{cor}(f,\,P+f_{<d-1})=\operatorname{bias}(f-(P+f_{<d-1}))=\operatorname{bias}(g+h-P)$
$=\operatorname{cor}(g+h,\,P)\geq 1/q^{\tilde{O}_{d}(r)}=\operatorname{bias}(\Delta^{d}f)^{\tilde{O}_{d}(1)}.$

### 30 Apr 2026 v2

@@ 6.2 Main proof
We now put everything together—in particular, Propositions 6.4, 4.5, 4.7 and Theorem 2.1—to prove Theorem 1.
@@
$\operatorname{rk}(h-\partial_{c}g)\leq\operatorname{rk}(\overline{h}-\partial_{c}\overline{g})\quad\text{ and }\quad\operatorname{rk}(g)\leq\operatorname{rk}(\overline{g}).$ (7)
To see this, abbreviate $\mathbf{x}^{t}=(x,\ldots,x)$ ($t$ times), and observe that $\partial_{c}\overline{g}(\mathbf{x}^{d})=\overline{\partial_{c}g}(\mathbf{x}^{d})=d!\partial_{c}g(x)$, using Lemma 5.2 and Fact 5.7. Moreover by Fact 5.7, we have $\overline{h}(\mathbf{x}^{d})=d!h(x)$ and $\overline{g}(\mathbf{x}^{d+1})=(d+1)!g(x)$. As $\operatorname{ch}(\mathbb{F})>d+1$, we thus have $g(x)=\overline{g}(\mathbf{x}^{d+1})/(d+1)!$ and $(h-\partial_{c}g)(x)=(\overline{h}-\partial_{c}\overline{g})(\mathbf{x}^{d})/d!$, proving (7) (since plugging in the diagonal in a reducible multilinear form gives a reducible form).
By Proposition 6.4, $\operatorname{bias}(\Delta^{d}f)\leq\operatorname{bias}(\overline{g})$, and $\operatorname{bias}(\Delta^{d}f)\leq\operatorname{bias}(\overline{h}-\partial_{c}\overline{g})$ for some direction vector $c$ which we henceforth fix. Put $\operatorname{bias}(\Delta^{d}f)=q^{-r}$ (if $\operatorname{bias}(\Delta^{d}f)=0$ there is nothing to prove), so that
@@
Now, combine (7), (8), and Theorem 2.1 (using $\operatorname{rk}(T)\leq\operatorname{PR}(T)$) to deduce that
@@
$\operatorname{rk}^{*}(g+h+\gamma)\leq\operatorname{rk}(g)+\operatorname{rk}(h-\partial_{c}g)\leq\tilde{O}_{d}(r)$
for some $\gamma\in\operatorname{Poly}_{d-1}$. Next, apply Proposition 4.7 with $k=d+1$ ($\geq 3$ by assumption) on $g+h+\gamma\in\operatorname{Poly}_{d+1}$. We deduce that there exists a polynomial $P$ with $\deg(P)\leq k-2=d-1$ such that
$\operatorname{cor}(g+h+\gamma,\,P)\geq q^{-2\operatorname{rk}^{*}(g+h+\gamma)}\geq q^{-\tilde{O}_{d}(r)}.$
To finish, write $f=g+h+f_{<d}$. Note that $P+f_{<d}-\gamma\in\operatorname{Poly}_{d-1}$; we claim that it significantly correlates with $f$. Indeed,
$\operatorname{cor}_{<d}(f)$ $\geq\operatorname{cor}(f,\,P+f_{<d}-\gamma)=\operatorname{bias}(f-(P+f_{<d}-\gamma))=\operatorname{bias}(g+h+\gamma-P)$
$=\operatorname{cor}(g+h+\gamma,\,P)\geq q^{-\tilde{O}_{d}(r)}=\operatorname{bias}(\Delta^{d}f)^{\tilde{O}_{d}(1)}.$